# An $\ell^{1}$-Plug-and-Play Approach for MPI Using a Zero Shot Denoiser with Evaluation on the 3D Open MPI Dataset

### v2

@@ Abstract
Approach: We validate the reconstruction parameters of the method on a hybrid dataset and compare it with the baseline Tikhonov, DIP and the previous PP-MPI, which is a plug-and-play method with denoiser trained on MPI-friendly data.
Main results: We offer a quantitative and qualitative evaluation of the zero-shot plug-and-play approach on the 3D Open MPI dataset. Moreover, we show the quality of the approach with different levels of preprocessing of the data.
Significance: The proposed method employs a zero-shot denoiser which has not been trained for the MPI task and therefore saves the cost for training. Moreover, it offers a method that can be potentially applied in future MPI contexts.
@@ 1 Introduction
Magnetic Particle Imaging (MPI) is an emerging medical imaging modality that images the distribution of superparamagnetic nanoparticles by exploiting their non-linear magnetization response to dynamic magnetic fields. The principles behind MPI have been introduced in the Nature paper by Gleich and Weizenecker [17]. Later, Weizenecker et al. [66] produced a video (3D+time) of a distribution of a clinically approved MRI contrast agent flowing through the beating heart of a mouse, thereby taking an important step towards medical applications of MPI. Further, Gleich, Weizenecker and their team won the European Inventor Award in 2016 for their contributions on MPI. Since then, interest in the technology has been on the rise both with view towards applications and with view towards improving the technology on the hardware side as well as on the algorithmic side.
@@
There are various medical applications such as cancer detection [54] and cancer imaging [15, 58, 69], stem cell tracing [12, 29, 42], inflammation tracing and lung perfusion imaging [71], drug delivery and monitoring [72], cardiovascular [6, 60, 62] and blood flow [16] imaging as well as brain injury detection [46]. MPI can be also used within multimodal imaging [3]. Further applications can be the tracking of medical instruments [25], as well as MPI-assisted stroke detection and brain monitoring based on recent human-sized MPI scanner [20]. Benefits of MPI are quantifyability and high sensitivity, and compared with other imaging modality such as CT [10], MRI [50], PET [59] and SPECT [40], its ability to provide higher spatial resolution in a shorter acquisition time and the absence of ionizing radiation or radioactive tracers [37]. Further details concerning comparison with existing imaging modalities and further examples of state-of-the-art applications can be found in [8, 67].
@@
In brief, the principle and the procedure behind MPI are as follows: (i) the specimen to be scanned is injected with a tracer containing the magnetic nanoparticles (MNs) and placed in the scanner; (ii) a dynamic magnetic field with a low-field volume (LFV) is applied and the LFV is spatially moved; (iii) the LFV acts as a sensitive region whose motion causes the particles to induce a voltage in the receive coils. The imaging task of MPI is the reconstruction of the spatial distribution of the particles in the tracer from the voltage signal acquired during a scan. There are currently two main types of approaches to reconstructing particle distributions: measurement-based approaches and model-based approaches. Measurement-based approaches typically acquire a system matrix: for each column of the matrix, one performs scanning of a delta probe located in a corresponding voxel of a chosen 3D voxel grid [35, 65, 48, 41]. More precisely, a 3D grid of the volume to be scanned inside the scanner’s bore is considered; a probe with a reference concentration of tracer is iteratively positioned and scanned at each voxel of the grid. This way the response of the scanning system to discrete delta impulses at each voxel position is collected and stored in the system matrix $A$. Assuming linearity, the acquired signal $f$ of the scan of a specimen is then obtained as a superpositon of the scans of the delta impulses. For reconstruction, one has to solve a corresponding system of equations $Au=f,$ where the symbol $u$ denotes the desired concentration; loosely speaking, one has to invert the system matrix for the measured signal $f$. Due to the presence of noise and the ill-conditioning of the system matrix (e.g., [38, 56]), regularization techniques are needed for the inversion as discussed later. The development of model-based reconstruction techniques is a direction of research aiming at avoiding or significantly reducing the calibration procedures needed for measuring the system matrix; for details, cf. for instance  [47, 51, 18, 19, 21, 45, 9]. In this work we consider the measurement approach and focus on regularized machine learning-based reconstruction in that scenario.
Related Work
@@
Due to the ill-conditioning of the system matrix and the related ill-posedness of the reconstruction problem in MPI [34, 45, 33], small perturbations in the data can lead to significant errors in the reconstruction; thus, regularization is needed [7]. Classical regularization approaches in MPI include Tikhonov regularization [66, 38] using conjugate gradients and the Kaczmarz method. To improve noise suppression capabilities and further features, structural priors based on the $\ell^{1}$-norm have been used within iterative schemes; e.g.[56]. For a broader overview we refer to [13].
@@
A particular class of machine learning-based approaches are Plug and Play (PnP) approaches. They are inspired by iterative schemes realizing energy minimization: building on a splitting scheme, they replace the denoiser implied by the slitting with an arbitrary one; instead of classical denoising, typically machine learning-based denoisers are employed [70]. PnP approaches have been used for a variety of imaging modalitites, e.g., [2, 55] as well as for computer vision tasks, e.g., [70, 57]. These approaches are conservative in the sense that machine learning only enters the scheme via iterative denoising and that well-known and well-understood components from iterative schemes and classical Tikhonov regularization are still integrated into the reconstruction scheme.
Most closely related to the present work is the paper [4] which also considers a plug and play prior for MPI reconstruction. Our approach differs in the underlying iterative scheme and the denoising method used, which is at the heart of a PnP approach, and in the employment of an additional $\ell^{1}$-prior. We further provide both quantitative and qualitative evaluation on the full 3D Open MPI data set [39]. A detailed discussion concerning the novelties compared to [4] can be found in the discussion in Section 4.1.
Contribution
In this work we consider a plug and play (PnP) approach for MPI reconstruction based on a generic (zero-shot) denoiser, a half quadratic splitting scheme plus an additional $\ell^{1}$-prior. Moreover, we validate the parameters on a hybrid dataset and evaluate the proposed methods on the publicly available Open MPI dataset. In detail, we make the following contributions:
@@
    In the proposed PnP method we employ a generic (zero-shot) image denoiser. This means the denoiser is not transfer learned on MPI or similar MRI data. (To apply the 2D image denoiser on a 3D volume, we use slicing.) The approach is inspired by the recent observations in Computer Vision that larger models pretrained on larger training sets perform well even when employed in off-distribution contexts. We further employ a shrinkage/$\ell^{1}$-prior which is motivated by the works [56, 32] which use a corresponding prior in an energy minimization framework in MPI reconstruction.
@@
    We validate the parameters on a hybrid 3D MPI dataset.
@@
    We quantitatively and qualitatively evaluate our approach on the full 3D Open MPI dataset [39] which presently represents the only publicly available real MPI dataset. In particular, we quantitatively and qualitatively compare our method with the DIP approach [14] using the preprocessing pipeline in [31], with the classical Tikhonov regularization and previous PnP approaches [4]. Moreover, we investigate the influence of the preprocessing on the reconstruction quality.
Both quantitative and qualitative evaluation of a PnP approach (including comparison with another machine learning approach) on the full 3D Open MPI dataset has not been done before. (We note that quantitative evaluation is possible on the Open MPI dataset since a reference CAD model is available in the dataset and serves the purpose of a ground truth [39]). Further, employing a zero-shot denoiser [70] (trained on a large set of natural images) in the context of MPI is new.
Outline
In Section 2 we develop the method: we describe the formal MPI reconstruction problem (Section 2.1), the preprocessing pipeline of the MPI data (Section 2.2), we present the proposed ZeroShot-PnP approach with and without the additional $\ell^{1}$-prior (Section 2.3), we consider parameter selection in Section 2.4 and describe the hybrid dataset created for the validation of the parameters in Section 2.5. In Section 3 we quantitatively and qualitatively evaluate the proposed ZeroShot-$\ell^{1}$-PnP algorithm on the OpenMPI Dataset: we compare with the DIP, Tikhonov regularization and previous Plug-and-Play approaches (Section 3.3 and 3.4), investigate the influence of different levels of preprocessing (Section 3.5) and consider a setup without preprocessing in Section 3.6. We conclude with Section 4 which contains a discussion and a conclusion part.
@@ 2.1 The Reconstruction Problem in MPI Using a System Matrix
The task of MPI reconstruction is to determine the concentration $u$ of iron oxide nanoparticles contained in an object. We briefly describe the basic principle of a field-free point (FFP) scanner; for details we refer to [37]. A specific static magnetic field is applied to saturate the nanoparticles everywhere in a field of view but in a small region. Such low-field region is a small neighborhood of a point called the Field-Free Point (FFP). A second dynamic magnetic field is applied to steer the FFP in the field of view. The movement of the FFP along a given trajectory constitutes a scan: the change of magnetization of the iron oxide particles due to the movement of the FFP induces a voltage in a set of receive coils which constitutes the measured data $f$. The relation between the concentration $u$ and the measured data $f$ is modeled as a linear mapping called the system function or system matrix $A$. In order to determine the system matrix/function both model and measurement-based approaches have been applied; cf. [48, 36] for instance. At present, the measurement-based approach is the most popularly used in real data scenarios: the system matrix is obtained by measuring the system responses to delta probes centered at a voxel. We consider MPI reconstruction in three dimesional space using real measurement data from the Open MPI dataset [39]. Hence, the reconstructed 3D image/volume $\tilde{u}$ lives in the space $X=\mathbb{R}^{n_{1}\times n_{2}\times n_{3}},$ where $n_{1}=n_{2}=n_{3}=19$ for the Open MPI data set. We denote the total number of voxels by $N=n_{1}n_{2}n_{3}=19^{3},$ and the number of measurements by $M$; here $M$ depends on the preprocessing applied to the MPI data (cf. Section 2.2). In particular, a scan in the Open MPI dataset contains the data coming from three different channels and the calibration data is Fourier transformed. This means that for each of the three channels the real and imaginary parts of the Fourier coefficients are considered and stacked, yielding a real-valued submatrix $A_{l}$, where $l=1,2,3$ is the channel index. Finally, the full matrix $A$ is formed by stacking the submatrices $A_{l}$ for each channel $l$ (for a detailed explanation we refer the interested reader to [31]). In order to avoid notation issues, we use the following conventions: we use the symbol $u\in\mathbb{R}^{N}$ to denote the (columnwise) vectorization of the spatial distribution $\tilde{u}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}.$ We then identify the system matrix with a matrix $\tilde{A}\in\mathbb{R}^{M\times N}$ such that $\tilde{A}\tilde{u}=Au$ for all volumes $\tilde{u}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$. We further identify the adjoint $A^{\ast}$ of $A$ with the matrix $\tilde{A}^{T}$. Using this notation, we have the relation
$f=Au+\eta=\tilde{A}\tilde{u}+\eta.$ (1)
@@
where $\tilde{A}\in\mathbb{R}^{M\times N}$ denotes the system matrix, $u$ denotes the discrete nanoparticle concentration, $f$ the measured data, and the symbol $\eta$ denotes the noise on the measurements. The reconstruction task consists of determining $u$ in (1) given measurements $f$ (and calibration system matrix $A$). This is challenging due to noise and the ill-conditioning of $A$, and requires regularization.
@@ 2.2 The Preprocessing of the Data
Bandpass Filtering and Frequency Selection
Due to analog filtering, the measurement data show a very high variance in the frequencies up to 75kHz. We theqrefore exclude frequencies lower than 80kHz. Further frequency selection avoiding frequencies with systematic strong noise or with low signal-to-noise ratio (SNR) has been considered in [31]. To reduce the number of variables at play and to avoid estimations of the SNR we consider no SNR-based thresholding in the preprocessing here.
Background Correction
The noisy measured phantom data $f$ is of the form $f=f_{\mathrm{cl.}}+b_{f}+\eta_{f}$, with clean signal $f_{\mathrm{cl.}}$ containing the information on $u,$ the background signal $b_{f}$ produced by the drive field and perturbations in the measurement chain, and noise $\eta_{f}$ (assuming it is additive in nature). Accordingly, we decompose the system matrix as $A=A_{\mathrm{cl}}+b_{A}+\eta_{A},$ with clean signals $A_{\mathrm{cl}},$ background $b_{A}$ and noise $\eta_{A}.$ To exclude the background signals $b_{f},$ $b_{A}$, background measurements with no phantom in the scanner are performed after the scan and during the calibration procedure. To account for the background signal to be dynamic, it is repeated after every $19$th delta probe scan (after running through the $x$-axis) and a corresponding convex combination is subtracted. Each phantom and delta scan is performed $1000$ times and the average is taken as input for the reconstruction. By linearity, the averaging and subtraction is in fact carried out in the Fourier domain. We refer to [31] for more details.
Whitening Estimation and Low Rank Approximation of the System Matrix
Following [14] we consider a whitening matrix $W$ obtained from the diagonal covariance matrix of multiple background scans, i.e., we consider the diagonal matrix $W\in\mathbb{R}^{M\times M}$ whose diagonal entries are $1/\sigma_{b}$ where $\sigma_{b}^{2}$ is the variance of the background scans provided. When whitening is applied, we consider the whitened system matrix $WA$ and the whitened data $Wf$ instead of $A,$ $f$.
@@
2.3 ZeroShot-PnP: The Proposed Plug-and-Play Approach with a Zero-Shot Denoiser
@@
$\hat{u}=\arg\min_{u}\lVert f-Au\rVert_{2}^{2}+\lambda\mathcal{R}(u),$ (2)
where $\mathcal{R}$ is a regularizer and $\lambda>0$ is a parameter controlling the strength of the regularization. In the context of MPI, adding positivity and sparsity constraints has turned out to be effective [56, 32]. Adding positivity and sparsity contrains results in a minimization problem of the form
$\hat{u}=\arg\min_{u}\lVert f-Au\rVert_{2}^{2}+\lambda\mathcal{R}(u)+\alpha\lVert u\rVert_{1}+\iota_{+}(u),$ (3)
where $\iota_{+}$ is the indicator function defined as $\iota_{+}(u)=0$ if the components $u_{i}$ of $u$ are all $u_{i}>0$ and $\iota_{+}(u)=+\infty$ otherwise.
The Iterative Scheme
Choosing auxiliary variables $u_{1}$, $u_{2}$ and $u_{3}$ we decouple the quantities in Equation (3), in particular, we choose the following equivalent constrained minimization problem in consensus form:
@@
$u_{3}^{k+1}$ $=\arg\min_{u_{3}}\frac{\alpha}{\mu_{k}}\lVert u_{3}\rVert_{1}+\frac{1}{2}\lVert u_{1}^{k+1}-u_{3}\rVert_{2}^{2}.$ (8)
At each iteration $k$, the term in Equation (6) is in charge of the regularized inversion of the system matrix whereas the term in Equation (7) performs regularized Gaussian denoising with positivity constraint. Finally, the term in Equation (8) is the proximal mapping of the $\ell^{1}$-norm, i.e., it results in a soft-thresholding of parameter $\alpha/\mu_{k}$.
The Denoising Step
The idea behind the PnP approach is to substitute the Gaussian denoising in Equation (7) by any other denoiser with positive output and w.r.t. noise level $\lambda/\mu_{k}$, yielding the following algorithm:
@@
The choice of a suitable denoiser in Equation (10) is key because the performance of the algorithm significantly depends on it. In Computer Vision, machine learning based 2D denoisers have gained a lot of interest in recent years. They are typically trained on large datasets and possess a higher generalization ability than smaller models.
The denoiser we employ in this work is the benchmark deep denoiser prior [70]. It is based on a deep CNN architecture (DRUNet) which combines a U-Net [49] and ResNet [28]. The denoiser is publicly available and trained on a large dataset that combines various datasets including BSD [11], Waterloo Exploration Database [44], DIV2K [1] and Flick2K [43]. We do not retrain the denoiser on MPI or MPI-related data, but employ it as a zero-shot denoiser (in the sense that it has not seen the domain before).

### v3

@@ Abstract
Approach: We validate the reconstruction parameters of the method on a hybrid dataset and compare it with the baseline Tikhonov, ART, DIP and the previous PP-MPI, which is a Plug-and-Play method with denoiser trained on MPI-friendly data.
Main results: We derive a Plug-and-Play reconstruction method based on a generic zero-shot denoiser. Addressing (hyper)parameter selection, we perform an extended parameter search on a hybrid validation dataset we produced and apply the derived parameters for reconstruction on the 3D Open MPI Dataset. We offer a quantitative and qualitative evaluation of the zero-shot Plug-and-Play approach on the 3D Open MPI dataset with the validated parameters. Moreover, we show the quality of the approach with different levels of preprocessing of the data.
Significance: The proposed method employs a zero-shot denoiser which has not been trained for the MPI reconstruction task and theqrefore saves the cost for training. Moreover, it offers a method that can be potentially applied in future MPI contexts.
@@ 1 Introduction
Magnetic Particle Imaging (MPI) is an emerging medical imaging modality that images the distribution of superparamagnetic nanoparticles by exploiting their non-linear magnetization response to dynamic magnetic fields. The principles behind MPI have been introduced in the Nature paper by Gleich and Weizenecker [18]. Later, Weizenecker et al. [73] produced a video (3D+time) of a distribution of a clinically approved MRI contrast agent flowing through the beating heart of a mouse, thereby taking an important step towards medical applications of MPI. Further, Gleich, Weizenecker and their team won the European Inventor Award in 2016 for their contributions to MPI. Since then, interest in the technology has been on the rise both with view towards applications and with view towards improving the technology on the hardware side as well as on the algorithmic side.
@@
There are various medical applications such as cancer detection [60] and cancer imaging [16, 64, 76], stem cell tracing [13, 32, 46], inflammation tracing and lung perfusion imaging [80], drug delivery and monitoring [81], cardiovascular [6, 66, 68] and blood flow [17] imaging, tracking of medical instruments [27] as well as brain injury detection [52] and MPI-assisted stroke detection and brain monitoring based on recent human-sized MPI scanners [21]. MPI can be also used within multimodal imaging [3]. Benefits of MPI are quantifyability and high sensitivity, and compared with other imaging modalities such as CT [10], MRI [56], PET [65] and SPECT [44], its ability to provide higher spatial resolution in a shorter acquisition time and the absence of ionizing radiation or radioactive tracers [40]. Further details concerning comparison with existing imaging modalities and further examples of state-of-the-art applications can be found in [8, 74].
@@
In brief, the principle and the procedure behind MPI are as follows: (i) the specimen to be scanned is injected with a tracer containing the magnetic nanoparticles (MNs) and placed in the scanner; (ii) a dynamic magnetic field with a field-free region (FFR), e.g., a field-free point, is applied and the FFR is spatially moved; (iii) the FFR acts as a sensitive region whose motion causes the particles to induce a voltage in the receive coils. The imaging task of MPI is the reconstruction of the spatial distribution of the particles from the voltage signal acquired during a scan. There are currently two main types of approaches to reconstructing particle distributions: measurement-based approaches and model-based approaches. Measurement-based approaches typically acquire a system matrix: for each column of the matrix, one performs scanning of a delta probe located in a corresponding voxel of a chosen 3D voxel grid [38, 72, 54, 45]. More precisely, a 3D grid of the volume to be scanned inside the scanner’s bore is considered; a probe with a reference concentration of tracer is iteratively positioned and scanned at each voxel of the grid. This way the response of the scanning system to discrete delta impulses at each voxel position is collected and stored in the system matrix $A$. Assuming linearity, the acquired signal $f$ of the scan of a specimen is then obtained as a superpositon of the scans of the delta impulses. For reconstruction, one has to solve a corresponding system of equations $Au=f,$ where the symbol $u$ denotes the desired concentration distribution; loosely speaking, one has to invert the system matrix for the measured signal $f$. Due to the presence of noise and the ill-conditioning of the system matrix (e.g., [42, 62]), regularization techniques are needed for the inversion as discussed later. The development of model-based reconstruction techniques is a direction of research aiming at avoiding or significantly reducing the calibration procedures needed to measure the system matrix; for details, cf. for instance [53, 57, 19, 20, 23, 50, 9, 49]. In this work, we consider the measurement approach and focus on regularized machine learning-based reconstruction in that scenario.
Related Work.
@@
Due to the ill-conditioning of the system matrix and the related ill-posedness of the reconstruction problem in MPI [37, 50, 36], small perturbations in the data can lead to significant errors in the reconstruction; thus, regularization is needed [7]. Classical regularization approaches in MPI include Tikhonov regularization [73, 42] using conjugate gradients and the Algebraic Reconstruction Technique (ART), also known as the Kaczmarz method. To improve noise suppression capabilities and enhance the reconstruction of features such as shape, edges, corners etc., structural priors based on the $\ell^{1}$-norm have been used within iterative schemes [62, 51], as well as for sparse recovery of system matrices [22]. For a broader overview we refer to [14].
@@
A particular class of machine learning-based approaches are Plug-and-Play (PnP) approaches first introduced by Venkatakrishnan et al. [69]. They are inspired by iterative schemes realizing energy minimization: from a Bayesian perspective, the reconstruction of an image $x$ from (noisy) measurements $y$ can be achieved through Maximum a Posterior Estimate (MAP), i.e., by maximizing the conditional probability $\mathbb{P}(x|y)$; this maximization problem is equivalent to the minimization of the an energy functional of the type $\ell(y,x)+\lambda\mathcal{R}(x)$ where $\ell(y,x)$ is the log-likelihood and $\mathcal{R}(x)$ is a prior. By considering the log-likelihood in terms of a variable $x_{1}$ and the prior in terms of a variable $x_{2}$ and adding the constraint $x_{1}=x_{2}$, it is possible to formulate an equivalent but constrained minimization problem [69], which can be solved with a chosen splitting scheme in an alternated fashion, alternating between a ridge regression problem and a Gaussian denoising problem; instead of classical denoising, typically machine learning-based denoisers are employed [77] in the Gaussian denoising subproblems of the splitting scheme. PnP approaches have been used for a variety of imaging modalitites, e.g., [2, 61] as well as for computer vision tasks, e.g., [77, 63]. These approaches are conservative in the sense that machine learning only enters the scheme via iterative denoising and that well-known and well-understood components from iterative schemes and classical Tikhonov regularization are still integrated into the reconstruction scheme.
Most closely related to the present work is the paper [4] which also considers a Plug-and-Play prior for MPI reconstruction. Our approach differs in the underlying iterative scheme and the denoising method used, which is at the heart of a PnP approach, and in the employment of an additional $\ell^{1}$-prior. We further provide both quantitative and qualitative evaluation on the full 3D Open MPI data set [43]. A detailed discussion concerning the novelties compared to [4] can be found in the discussion in Section 4.1.
Contribution.
In this work we consider a PnP approach for MPI reconstruction based on a generic (zero-shot) denoiser, a half quadratic splitting scheme plus an additional $\ell^{1}$-prior. Moreover, we validate the parameters on a hybrid dataset and evaluate the proposed methods on the publicly available Open MPI dataset. In detail, we make the following contributions:
@@
    In the proposed PnP method we employ a generic (zero-shot) image denoiser. This means the denoiser is not transfer-learned on MPI or similar MRI data. (To apply the 2D image denoiser on a 3D volume, we use slicing.) We further employ a shrinkage/$\ell^{1}$-prior which is motivated by the works [62, 35, 51, 22] which use a corresponding prior in an energy minimization framework in MPI reconstruction.
@@
    We create a hybrid 3D MPI dataset by applying the system matrix in the Open MPI Dataset [43] to simulated phantoms as in [41, 79, 26]; we use this dataset to choose the reconstruction parameters that we later apply to real MPI data and hence provide a way to fix reconstruction parameters for real MPI data. Avoiding fine tuning of the hyperparameters on the data is beneficial in real world applications, as the ground truth is not available. In particular, we show here that the ZeroShot-$\ell^{1}$-PnP and ZeroShot-PnP algorithm obtain competitive results without tuning of the hyperparmeters on the data and with one common set of hyperparameters.
@@
    We quantitatively and qualitatively evaluate our approach on the full 3D Open MPI dataset [43] which presently represents the only publicly available real 3D MPI dataset. In particular, we quantitatively and qualitatively compare our method with the DIP approach [15] using the preprocessing pipeline in [34], with the classical Tikhonov regularization, the ART method [34, 42], and the previous PnP approach [4]. Moreover, we investigate the influence of the preprocessing on the reconstruction quality.
Both quantitative and qualitative evaluation of a PnP approach (including comparison with another machine learning approach) on the full 3D Open MPI dataset has not been done before. (We note that quantitative evaluation is possible on the Open MPI dataset since a reference CAD model is available in the dataset and serves the purpose of a ground truth [43]; we remark however, that such CAD ground truths are used exclusively for quantitative evaluation of the results and not for the choice of reconstruction parameters, which we perform on the hybrid dataset.) Furthermore, employing a zero-shot denoiser [77] (trained on a large set of natural images) in the context of MPI is new.
Outline.
In Section 2 we develop the method: we describe the formal MPI reconstruction problem (Section 2.1), the preprocessing pipeline of the MPI data (Section 2.2), we present the proposed ZeroShot-PnP approach with and without the additional $\ell^{1}$-prior (Section 2.3), we consider parameter selection in Section 2.4 and describe the hybrid dataset created for the validation of the parameters in Section 2.5. In Section 3 we quantitatively and qualitatively evaluate the proposed ZeroShot-$\ell^{1}$-PnP algorithm on the OpenMPI Dataset: we compare with the DIP, Tikhonov regularization, the ART method and the previous Plug-and-Play approach PP-MPI (Section 3.3 and 3.4), investigate the influence of different levels of preprocessing (Section 3.5) and consider a setup with most basic preprocessing consisting only of background removal and suppressing of excitation crosstalk in Section 3.6. We conclude with Section 4 which contains a discussion and a conclusion part.
@@ 2.1 The Reconstruction Problem in MPI Using a System Matrix
The task of MPI reconstruction is to determine the concentration $u$ of iron oxide nanoparticles contained in an object given the measured signal $f$ during a scan. The relation between the concentration $u$ and the measured data $f$ is modeled as a linear mapping called the system function or system matrix $A$. In order to determine the system matrix/function both model and measurement-based approaches have been applied; cf. [54, 39] for instance. At present, the measurement-based approach is the most popularly used in real data scenarios. We consider MPI reconstruction in three dimesional space using real measurement data from the Open MPI dataset [43]. Hence, the reconstructed 3D concentration $u_{\mathrm{rec}}$ lives in the space $X=\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}\times\mathbb{R}^{n_{3}},$ where $n_{1}=n_{2}=n_{3}=19$ for the Open MPI data set. We denote the total number of voxels by $N=n_{1}n_{2}n_{3}=19^{3},$ and the number of frequency components by $M$ (possibly upon transformation of the data if provided in time domain into the Fourier domain); here $M$ depends on the preprocessing applied to the MPI data (cf. Section 2.2). In particular, a scan in the Open MPI dataset contains the data coming from three different channels, i.e., from two couples of perpendicular receiving coils coaxial to the y and z dimensions, a solenoid in the x direction, and the calibration data is Fourier transformed if provided in time domain. This means that for each of the three channels the real and imaginary parts of the Fourier coefficients are considered and stacked, yielding a real-valued submatrix $A_{l}$, where $l=1,2,3$ is the channel index. Finally, the full matrix $A$ is formed by stacking the submatrices $A_{l}$ for each channel $l$ (for a detailed explanation we refer the interested reader to [34]). In order to avoid notational issues, we use the following conventions: $u_{\mathrm{GT}}\in\mathbb{R}^{N}$ denotes the (columnwise) vectorization of the ground truth spatial distribution, we will use a tilde to denote the 3D or 2D reshaping of $u_{\mathrm{GT}}$, i.e., $\tilde{u}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$. We identify the adjoint $A^{\ast}$ of $A$ with the matrix $A^{T}$. Using this notation, we have the relation
$f=Au_{\mathrm{GT}}+\eta,$ (1)
@@
where $f$ is the measured data, and the symbol $\eta$ denotes the noise on the measurements which affects both $f$ and each column of the system matrix $A$. The reconstruction task consists of determining a distribution $u_{\mathrm{rec}}$ approximating $u_{\mathrm{GT}}$ from (1) given measurements $f$ (and calibration system matrix $A$). This is challenging due to noise and the ill-conditioning of $A$, and requires regularization.
@@ 2.2 The Preprocessing of the Data
Bandpass Filtering and Frequency Selection.
During the collection of scan data, an analog filter is in place to suppress excitation crosstalk; consequently, the Signal-to-Noise ratio of frequencies below about 80 kHz is reduced [54]. For this reason, we always exclude frequencies lower than the 80 kHz threshold. Further frequency selection avoiding frequencies with systematic strong noise or with low signal-to-noise ratio (SNR) has been considered in [34] and are part of the possible preprocessing steps. To reduce the number of variables at play and to avoid estimations of the SNR we consider no SNR-based thresholding in the preprocessing here.
Background Correction.
The noisy measured phantom data $f$ is of the form $f=f_{\mathrm{cl.}}+b_{f}+\eta_{f}$, with clean signal $f_{\mathrm{cl.}}$ containing the information on $u,$ the background signal $b_{f}$ produced by the drive field and perturbations in the measurement chain, and noise $\eta_{f}$ (assuming it is additive in nature). Accordingly, we decompose the system matrix as $A=A_{\mathrm{cl}}+b_{A}+\eta_{A},$ with clean signals $A_{\mathrm{cl}},$ background $b_{A}$ and noise $\eta_{A}.$ To exclude the background signals $b_{f},$ $b_{A}$, background measurements with no phantom in the scanner are performed after the scan and during the calibration procedure. To account for the background signal to be dynamic, in the calibration procedure a background scan is repeated after every $19$th delta probe scan (after running through the $x$-axis) and a corresponding convex combination is subtracted. Each phantom and delta scan is performed $1000$ times and the average is taken as input for the reconstruction. By linearity, the averaging and subtraction is in fact carried out in the Fourier domain. We refer to [34] for more details.
Whitening Estimation and Low Rank Approximation of the System Matrix.
Following [15], we consider a whitening matrix $W$ obtained from the diagonal covariance matrix of multiple background scans, i.e., we consider the diagonal matrix $W\in\mathbb{R}^{M\times M}$ whose diagonal entries are $1/\sigma_{b}$ where $\sigma_{b}^{2}$ is the variance of the background scans provided. When whitening is applied, we consider the whitened system matrix $WA$ and the whitened data $Wf$ instead of $A$ and $f$.
@@
2.3 ZeroShot-PnP: The Proposed Plug-and-Play Approach with a Zero-Shot Denoiser.
@@
$u_{\mathrm{rec}}=\arg\min_{u}\lVert f-Au\rVert_{2}^{2}+\lambda\mathcal{R}(u),$ (2)
where $\mathcal{R}$ is a regularizer and $\lambda>0$ is a parameter controlling the strength of the regularization. In the context of MPI, adding positivity constraints is motivated by the fact that the solutions $u_{\mathrm{rec}}$ in Eq. (2) represent particle distribution, and additional $\ell^{1}$-priors have turned out to be effective in MPI thanks to their sparsifying effect on the solution [62, 35, 51, 31]. Adding positivity and sparsity constraints results in a minimization problem of the form
$u_{\mathrm{rec}}=\arg\min_{u}\lVert f-Au\rVert_{2}^{2}+\lambda\mathcal{R}(u)+\alpha\lVert u\rVert_{1}+\iota_{+}(u),$ (3)
where $\iota_{+}$ is the indicator function defined as $\iota_{+}(u)=0$ if the components $u_{i}$ of $u$ are all $u_{i}>0$ and $\iota_{+}(u)=+\infty$ otherwise, and $\alpha>0$ is an additional regularization parameter controlling strength of the $\ell^{1}$-prior.
The Iterative Scheme.
Choosing auxiliary variables $u_{1}$, $u_{2}$ and $u_{3}$, we decouple the quantities in Equation (3), in particular, we choose the following equivalent constrained minimization problem in consensus form:
@@
$u_{3}^{k+1}$ $=\arg\min_{u_{3}}\frac{\alpha}{\mu_{k}}\lVert u_{3}\rVert_{1}+\frac{1}{2}\lVert u_{1}^{k+1}-u_{3}\rVert_{2}^{2}\quad.$ (8)
We follow [77] and employ a continuation strategy, i.e., a varying $\mu_{k}$ in Equation (6); using parameters varying in each iteration have already been employed in non MPI-related contexts, for example in [11]. The continuation strategy is one of the differences with the previous PP-MPI approach [4]. Concerning the HQS splitting, it is possible to assign a helpful interpretation to the subproblems: at each iteration $k$, the term in Equation (6) is in charge of the regularized inversion of the system matrix whereas the term in Equation (7) performs regularized Gaussian denoising with positivity constraint; if we define the combined parameter $\sigma_{k+1}^{2}\coloneq\lambda/\mu_{k}$, because Equation (7) is equivalent in form to a Gaussian denoising problem, the parameter $\sigma_{k+1}^{2}$ can be interpreted as the variance of the noise of $u_{1}^{k+1}$ and $\sigma_{k+1}=\sqrt{\lambda/\mu_{k}}$ as its noise level. Finally, the term in Equation (8) is the proximal mapping of the $\ell^{1}$-norm, i.e., it results in a soft-thresholding of parameter $\alpha/\mu_{k}$.
The Denoising Step.
The idea behind the PnP approach is to substitute the denoising problem in Equation (7) by any other denoiser with positive output and w.r.t. noise level $\sqrt{\lambda/\mu_{k}}$, yielding the following algorithm:
@@
The choice of a suitable denoiser in Equation (10) is key because the performance of the algorithm depends on it. In Computer Vision, machine learning based 2D denoisers have gained a lot of interest in recent years. They are typically trained on large datasets and possess a higher generalization ability than smaller models [77].
The denoiser we employ in this work is the benchmark deep denoiser prior [77]. It is based on a deep CNN architecture (DRUNet) which combines a U-Net [55] and ResNet [30]. The denoiser is publicly available and trained on a large dataset that combines various datasets including BSD [12], Waterloo Exploration Database [48], DIV2K [1] and Flick2K [47]. During training of the deep denoiser prior, the authors in [77] considered 16 patches of size $128\times 128$ randomly cropped out of the training dataset (image in the range $[0,255]$) and added to them additive white Gaussian noise with noise level $\sigma$ randomly chosen from $[0,50]$, to account for big variations in the noise level. In particular, the deep denoiser prior takes as input the noisy image and a noise level map, which is a uniform map filled with the value $\sigma$ and with the same size as the image. In the PP-MPI prior in [4], the denoiser does not take a noise level map as input; here lies the main difference between the PP-MPI approach and our ZeroShot-PnP approach: the possibility of inputting a noise level map in the deep denoiser prior allows to couple the parameter $\mu_{k}$ in Eq. (9) with the noise level $\sqrt{\lambda/\mu_{k}}$ in Eq. (10) and design an automatic parameter selection strategy that we will describe in Section 2.4. Moreover, we do not retrain the deep denoiser prior on MPI or MPI-related data, but employ it as a zero-shot denoiser (in the sense that it has not seen the domain before).